### v1

@@ Abstract
Global food systems must deliver nutritious and sustainable foods while sharply reducing environmental impact. Yet, food innovation remains slow, empirical, and fragmented. Artificial intelligence (AI) now offers a transformative path with the potential to link molecular composition to functional performance, bridge chemical structure to sensory outcomes, and accelerate cross-disciplinary innovation across the entire production pipeline. Here we outline AI for Food as an emerging discipline that integrates ingredient design, formulation development, fermentation and production, texture analysis, sensory properties, manufacturing, and recipe generation. Early successes demonstrate how AI can predict protein performance, map molecules to flavor, and tailor consumer experiences. But significant challenges remain: lack of standardization, scarce multimodal data, cultural and nutritional diversity, and low consumer confidence. We propose three priorities to unlock the field: treating food as a programmable biomaterial, building self-driving laboratories for automated discovery, and developing deep reasoning models that integrate sustainability and human health. By embedding AI responsibly into the food innovation cycle, we can accelerate the transition to sustainable protein systems and chart a predictive, design-driven science of food for our own health and the health of our planet.
keywords:
artificial intelligence, machine learning, human health, planetary health, alternative protein
Motivation
@@
Food is one of humanity’s most consequential technologies for human health, planetary sustainability, and cultural identity [1]; yet, research and development for food remains slow, fragmented, and driven by trial-and-error rather than rational design. Radical, urgent, and creative food solutions are needed to combat the current climate crisis [2]. Strikingly, the food sector contributes 35% of the global green house emissions across farm land, livestock, and production [3] and contributes heavily to biodiversity loss and public health challenges [4]. To meet climate goals and feed a growing population, we need alternative sources of protein that are delicious, affordable, scalable, and nutritionally compelling [5]. Today, however, the scientific bottlenecks in texture, flavor, scalability, and cost pose barriers to adoption [6]. Artificial intelligence (AI) has transformed the scientific landscape, entirely reshaping domains from protein design and drug discovery to materials engineering [7]. AI is now poised to revitalize how we grow and make food – from optimizing plant-based meat formulations to transforming menus with climate-conscious recipes [8]. Enabling protein diversification will have profound implications, and AI can substantially accelerate scientific progress if applied responsibly and towards the correct challenges [9]. But this acceleration is not guaranteed. If deployed in a scattershot way, AI risks generating low-quality outputs, increasing environmental and resource burdens, or undermining consumer trust, aspects that are especially sensitive in food [10]. The goal is not to deploy AI indiscriminately. Instead, we need to partner with AI where it measurably reduces time, cost, and uncertainty for problems that matter [11]. Responsible application–focused on the right challenges–is essential. In this perspective, we identify areas across the food innovation where we can constructively engage with AI tools, we provide background on the current state of the art, outline new opportunities, and discuss the challenges facing AI to accelerate the future of food.
Figure 1: Historical milestones in AI for Food. Enabling discoveries in artificial intelligence (AI) for food include the birth of artificial intelligence, the rise of machine learning (ML) and neural networks (NN), the era of the human genome project, and the time of big data and cloud computing laid the foundation for AI for food. Since the beginning of the 21st century, developments in computational chemistry, deep learning, generative models, and fermentation have inspired the use of these technologies for food innovation. The current decade is witnessing the first use of artificial intelligence for personalized nutrition, the application of digital twins in food manufacturing, and the use of foundation models and large language models (LLM) for food, which are collectively fueling the emergence of AI for Food as its own scientific discipline.
@@
Historical Milestones in AI for Food. The intellectual and technological trajectory that has shaped to the emergence of AI for Food as a discipline spans more than seven decades, weaving together advances in artificial intelligence, food science, and biotechnology. From the birth of symbolic AI in the 1950s [12] and the rise of machine learning and neural networks in the 1980s–90s [13], to the deep learning breakthroughs of the 2010s [14], and the advent of generative models [15] and foundation models [16] in the 2020s, AI has progressively expanded its ability to model, design, and predict complex systems. In parallel, developments in genomics [17], computational chemistry [18], fermentation technologies and personalized nutrition [19] created unprecedented data resources and biological insights directly relevant to food design. Their convergence has given rise to a new field in which AI is now central to ingredient discovery, recipe formulation, manufacturing optimization, and consumer personalization [11]. Figure 1 summarizes this evolution and highlights twelve pivotal milestones that trace the path from early AI to today’s emerging discipline of AI for Food.
Figure 2: The promise of AI for Food. AI is already transforming food innovation. Representative AI-powered foods include meat alternatives in pink: burgers, steaks, deli slices, and sausages; dairy alternatives in blue: milk, cheese, ice cream, and yogurt; baked goods and snacks in orange: cookies, bread, chips, and protein bars; and condiments and ingredients in green: mayonnaise, ketchup, soups, and salad dressings. Together, these categories highlight AI’s role in creating healthier, tastier, and more sustainable foods.
@@
The Promise of AI for Food. Designing food materials poses unique challenges: its properties emerge from multiscale interactions among proteins, carbohydrates, lipids, and their processing conditions [20]. Yet, AI has moved food innovation from slow trial-and-error to data-driven discovery. Figure 2 illustrates four successful AI-powered food categories ranging from meat and dairy alternatives to snacks and condiments. Modern AI models now link chemical composition to sensory outcomes, enabling targeted formulation and rapid iteration. Machine-learning frameworks have predicted consumer appreciation directly from chemical–sensory panels in beverages [21], while graph neural networks have mapped molecular structure to odor quality, offering generalizable approaches to flavor design [22]. Advances in protein modeling and molecular machine learning accelerate the discovery and engineering of functional ingredients [23, 24], while synthetic-biology–enabled precision fermentation provides scalable routes to novel proteins and textures [25]. Importantly, these innovations begin to address long-standing challenges: creating plant-based burgers and steaks with the fibrous texture of animal muscle, designing plant-based milk and cheese with the emulsification and meltability of dairy proteins, and producing condiments like mayonnaise without egg-derived stabilizers. Crucially, start-ups including NotCo, Perfect Day, Meati, and Climax Foods have translated these methods into marketable foods [26]. This shift from speculative promise to market reality exemplifies how AI now functions as a powerful engine of culinary innovation and sustainability.
The AI-Powered Production Cycle. Our proposed development model for plant-based, fermentation-derived, or cultivated products follows an AI-powered cycle that moves through ingredient design, formulation development, fermentation and production, texture analysis, sensory properties, manufacturing, and recipe generation.
@@
Figure 3 illustrates this iterative cycle, in which each step presents an opportunity for AI to link experimental data with human perception and create a faster and more tightly connected development process. Traditional trial-and-error methods cannot scale to the complexity of modern food design, but AI can unlock design principles that open entirely new pathways for food innovation. By closing the loop between biology, food science, engineering, and consumer perception, this cycle enables continuous refinement from raw ingredients to commercial launch. In the following sections, we highlight how state-of-the-art approaches, from generative models to digital twins, can drive innovation across every step of this process.
AI for Ingredients
@@
Motivation. Novel food ingredients must balance sustainability, nutrition, sensory appeal, and functionality. Functionality refers to techno-functional traits such as water-holding, emulsification, gelation, viscosity, and stability, which shape processing and final structure. Many novel ingredients are designed to replicate animal-derived functionalities, for example, foaming capacity in egg replacers, yet achieving comparable performance remains a significant challenge. Structure-function relationships, well established in fields such as medical research, link molecular architecture to activity. We propose extending this concept to food innovation, whereby the macromolecular features of proteins, lipids, and fibers encode functionality of the ingredient as a whole. Can AI bridge this scale gap by translating molecular features and interactions into predictions of techno-functional traits, enabling a functionality-first design–build–test–learn cycle?
Background and State of the Art. The success of animal-free ingredients depends heavily on their functionality and the extent to which formulation and consumer requirements can be met [27, 29, 28]. These traits arise from molecular features, e.g., sequence and conformation; from environmental factors, e.g., pH, ionic strength, and temperature; and from interactions with fats, fibers, and solutes [28]. Proteins, lipids, and fibres provide the core techno-functional properties, with proteins offering the greatest versatility. Wheat gluten and soy-based proteins currently dominate due to their animal-like functionality. Gluten’s disulfide-linked networks, for example, convert mesoscopic structures into fibrous textures under extrusion and shear [27]. To move beyond these staples, attention is turning to legumes, oilseeds, grains, algae, and leaves [30, 31]. In addition to input sources, processing strongly influences functionality: dry fractionation yields protein-enriched flours with largely native structures, while wet fractionation with isoelectric precipitation produces concentrates and isolates that shape solubility, gelation, and dispersibility [30, 32]. Lipids and fibers provide complementary functionality. Oleosome emulsions enhance interfacial stability [33], while fibres modulate viscosity, emulsions, and lubrication, offering routes to mimic animal fat [34, 35]. However, the inherent heterogeneity of plant-derived inputs complicates predictable performance. To address this, ingredients are often modified chemically, physically, or enzymatically. Increasingly, a shift is underway toward functionality-enriched fractions rather than purified isolates, aiming for more consistent and cost-effective performance [30, 28, 31]. In parallel, bottom-up strategies such as colloid design, tuning droplet size, interfacial layering, and emulsion/foam-gel networks, enable direct control of macroscale texture and stability [36].
@@
Applications and Opportunities. AI and informatics now enable functionality-driven ingredient design by linking molecular and mesoscopic features to macroscopic outcomes under real processing constraints. Classification models support ingredient discovery and repurposing; for example, a quantitative structure–activity relationship model identified glycyrrhizin as a natural emulsifier, later validated experimentally [37]. Regression models enable rapid quality control by predicting pectin viscosity from simple parameters, lowering costs and accelerating formulation adjustments [38]. Functionality coupled to sensory design is increasingly supported by deep learning [39]. One study optimized starch-based gels for saltiness perception and gel strength by integrating e-tongue data with oral-processing measurements [40]. These gels are already used to reduce fat and sodium in animal-derived products and offer opportunities for animal-free alternatives. Explainable AI adds interpretability, for instance by showing how ionic conditions shape sesame protein rheology and thereby affect shear stress and viscosity [41]. Multi-objective optimization frameworks now predict solubility, gel strength, and emulsification capacity across diverse crops, providing scalable strategies for formulation [34]. Together, these advances enable closed-loop design–build–test–learn cycles, an iterative process of designing, constructing, testing, and refining ingredients. High-throughput screening, active learning, and digital twins operationalize AI models, while state-of-the-art methods combine latent-space learning such as autoencoders, conditional generation, and constraint-aware optimization to target multiple criteria and bridge lab-to-plate gaps.
@@
Challenges and Open Questions. Ingredient design still struggles with consistent performance across diverse food matrices. Several major bottlenecks remain: first, model transferability poses a major challenge because AI models trained on specific ingredients or conditions often fail to generalize across new food matrices. Second, nonlinear system behavior complicates prediction, as multicomponent food systems exhibit emergent interactions that require hybrid approaches combining physics-informed machine learning with empirical data. Third, texture and sensory preferences vary across cultures and regions, which means AI models must be trained on culturally diverse datasets to ensure consumer relevance. Fourth, interpretability and scalability remain essential, since models need to transparently link molecular features such as sequence, charge, or hydrophobicity to techno-functional performance. Resolving these challenges requires open, collaborative tools for functionality-first ingredient design, with the ultimate vision of an AI-powered discovery ecosystem that predicts techno-functional performance from molecular features.
AI for Formulations
@@
Motivation. Ingredient functionality sets the foundation, but effective food design depends on how those ingredients interact in complete systems; formulation brings these components together. Meeting the dual mandate of sustainability and sensory fidelity in alternative proteins requires moving beyond slow, trial-and-error formulation development. AI offers a compelling approach to uncover hidden patterns that connect ingredients, formulations, production, texture, and sensory properties, and propose candidate formulations that are predicted to meet nutritional, functional, or sensorial targets [42, 43, 44, 45]. Today, AI practitioners can draw on a broad collection of public resources including Recipe1M+ [46], FlavorGraph [47], FlavorDB [48], USDA FoodData Central, and FAO/INFOODS, as well as domain-specific databases such as Phenol-Explorer, LipidBank, and the Volatile Compounds in Food database. Inputs for AI models include molecular descriptors, nutritional and functional properties, recipe structures, and process sequences; outputs are complete candidate formulations aligned with target criteria. Figure 4 illustrates state-of-the-art methods that combine latent-space learning through autoencoders with conditional generation, constraint-aware optimization, and expert-in-the-loop refinement. Given the diversity of raw material information sources needed, can AI capture the complex, non-linear interactions between ingredients, processes, and perception–even under sparse and noisy data–and unlock a new generation of food formulations?
Figure 4: AI for formulations. An iterative workflow to discover new formulations encodes target profiles such as sensory properties, cost, and nutritional profiles; generates formulation sets with deep learning; predicts properties and scores; and integrates laboratory prototyping and sensory panels in continuous feedback loops.
@@
Background and State of the Art. Computational gastronomy has long shown that recipes and flavor chemistry form structured networks that can be mined for predictive patterns. The Flavor Network formalized ingredient compatibility via shared volatiles [49], while FlavorGraph integrated one million recipes with $\sim$1,500 flavor molecules to generate graph embeddings for pairing prediction and ingredient clustering [47]. These studies demonstrated that high-dimensional, multimodal food representations can capture the complexity of culinary design. In alternative proteins, industrial disclosures from NotCo patents [42, 43, 44, 45] illustrate hybrid generative optimization pipelines. These systems embed ingredients and recipes in latent spaces using autoencoders and recurrent networks, condition sampling on feature vectors of target product profiles, refine candidates with constraint solvers such as sparse regression, and iteratively incorporate expert sensory feedback. This architecture mirrors multimodal generative trends in other fields and enables analogue discovery. It can identify non-obvious ingredient substitutions while maintaining process feasibility. Other companies, including Climax Foods and Eat Just, deploy similar proprietary approaches, while academic research continues to build on Recipe1M+ [46], FlavorGraph [47], and FlavorDB [48]. Together, these efforts highlight a state of the art that blends representation learning with constraint-aware generation to accelerate formulation discovery.
@@
Applications and Opportunities. Generative AI opens new opportunities across protein diversification: It can help design plant-based cheeses that melt and stretch by identifying plant proteins with casein-like properties, since casein is the abundant milk protein that gives animal cheeses their characteristic structure and meltability. It can discover novel ingredient blends to replicate meat textures or pinpoint volatile combinations that evoke familiar flavors. It can brainstorm ideas for improving plant-based meat products in response to sensory panel feedback, and can even match an expert food scientist in predicting sensory panel rankings [98]. Large-scale virtual screenings allow researchers to focus on a handful of promising prototypes rather than having to rely on hundreds of trials, which shortens the production cycle. In the near future, these tools could tailor formulations to individual nutritional needs or taste preferences and unlock the full potential of personalized food design.
@@
Challenges and Open Questions. Significant challenges remain before AI can transform formulation discovery at scale. High-quality sensory datasets are limited, constraining generalization across different food types. Even with detailed chemical and nutritional data, predicting human perception remains an open problem. Modeling ingredient performance under varying heat, shear, or storage conditions is incomplete. Cost, supply chain, allergen control, and clean-label requirements must be built into models as hard constraints. Trust among regulators, industry, and consumers will depend on transparent, auditable design processes. The near-term path is hybrid, where AI proposes and humans refine, toward transparent formulation platforms that integrate constraints and predict performance at scale.
AI for Fermentation & Production
Motivation. While molecular and formulation-level design establishes the building blocks of alternative proteins, realizing their potential requires equally sophisticated control at the cellular and microbial scale. Fermentation is central to protein diversification and sustainable food production, yet its success depends on tightly coupled factors such as strain stability, metabolic fluxes, nutrient composition, and bioreactor dynamics including pH, oxygen, and agitation. The integration of domain knowledge with historical and real-time data through AI-driven strategies holds the potential to overcome technical bottlenecks in optimizing yields with reduced resources. Still, fermentation today often depends on trial and error because datasets across microbial strain engineering, metabolic responses, and upscaling are complex, fragmented, and difficult to integrate. Can AI fuse these diverse datasets with mechanistic insight to enable predictive, scalable, and cost-effective fermentation and production strategies for alternative proteins?
Figure 5: AI for fermentation and production. Artificial intelligence and large language models can accelerate the fermentation and production cycle during microbial strain selection; data collection; process optimization; real time monitoring and adaptive control; upscaling of prediction and simulation; and product profiling and customization.
@@
Background and State of the Art. Optimizing fermentation production requires a deep understanding of bioprocesses to maximize yield and efficiency while maintaining homogeneity, sterility, and scalability. Kinetic bioprocess models remain widely used but are constrained by steady-state assumptions, and metabolic flux models for engineered strains often face computational bottlenecks. Experimental optimization through orthogonal design [50] and response surface methodology [51] remains the industry standard, while multiscale feature engineering is still limited [52]. Machine learning is valuable for modeling nonlinear fermentation dynamics [53] and monitoring anaerobic digestion [54], but shallow models lack predictive power. Inline monitoring tools such as biosensors or in situ microscopy remain expensive and difficult to scale [56], and advanced control strategies rarely transition into industrial practice [55]. AI applications to sustainable protein fermentation–including strain engineering, process optimization, and real-time monitoring–are emerging [57], and reinforcement learning has been explored for co-culture optimization [58]. However, large-scale adoption of AI, and especially LLMs, remains limited.
@@
Applications and Opportunities. AI can shift fermentation and production from reactive optimization to predictive management, with customized strains and self-optimizing processes. Machine learning models can optimize strain design, nutrient feeds, and bioreactor control, and deep learning approaches can capture non-linear interactions better than traditional regression [59]. Automation and AI-enhanced imaging [60], soft sensors [61], and spectroscopy [57] can improve monitoring and adaptive control. Algorithms such as support vector machines [62] and fuzzy inference systems [63] support robust predictions, while evolutionary optimization [64] aids process tuning. Hybrid mechanistic-AI models [65] and explainable AI frameworks [66] can improve transparency and decision-making. Training LLMs on unstructured bioprocess data–including text, images, and sensor logs–could support knowledge retrieval, troubleshooting, and experimental design and bridging biology, chemistry, engineering and design through actionable insights, like predicting protein structures or cell morphology [67]. Figure 5 highlights the impact AI and large language models can have during the fermentation and production cycle. Collectively, these advances open opportunities for guided modeling, smart bioreactors, anomaly detection, forecasting, and multi-objective optimization across cost, nutrition, and sensory outcomes.
@@
Challenges and Open Questions. To realize this potential, the field must overcome the key barriers of data scarcity, data integration, and model transferability. Generating large experimental datasets remains labor intensive, though scaled-down bioreactors and small-data algorithms offer new possibilities. Robust biosecurity and ethical frameworks will be critical for LLM deployment. Integration with genome-scale metabolic models and digital twins of bioreactor performance could create predictive platforms for design and scale-up. The ultimate goal is an automated, multimodal bioprocessing platform that integrates diverse models to cut time and cost while ensuring reliable, explainable, and human-aligned decision-making.
AI for Texture

### v2

@@ Abstract
Global food systems must deliver nutritious, sustainable foods while sharply reducing environmental impact. Yet, food innovation remains slow, empirical, and fragmented. Artificial intelligence (AI) offers a transformative path to link molecular composition to functional performance, connect chemical structure to sensory outcomes, and accelerate cross-disciplinary innovation across the production pipeline. While broadly applicable to food systems, we focus on sustainable proteins–plant-based, fermentation-derived, and cultivated–as a high-impact testbed for AI-driven closed-loop design. We review the applications, opportunities, and challenges of AI for Food as an emerging discipline that integrates ingredient design, formulation development, fermentation and production, texture analysis, sensory science, manufacturing, and recipe generation. We identify four priorities: advancing scientific machine learning with embedded domain priors, treating food as a programmable biomaterial, building self-driving laboratories for automated discovery, and developing deep reasoning models that integrate nutrition and sustainability. Integrating AI responsibly into the food innovation cycle can accelerate the transition to sustainable food systems and establish a predictive, design-driven science of food for human and planetary health.
                 Editor’s summary
Artificial intelligence is transforming how new foods are designed, produced, and experienced. While AI has the potential to link molecular composition to functional and sensory performance and accelerate food innovation, challenges related to data quality, transparency, and governance remain. Addressing these requires interdisciplinary collaboration, open and inclusive data ecosystems, and AI systems that augment rather than replace human expertise.
Figure 1: Historical milestones in AI for Food. The birth of AI in the 1950s–1960s, marked by Alan Turing’s concepts of computing machinery and intelligence, established the foundations of machine intelligence. The rise of machine learning and neural networks in the 1980s–1990s, driven by backpropagation and statistical learning, shifted AI from rule-based systems to data-driven approaches. In parallel, the human genome project and the omics revolution in the 1990s–2000s, enabled by high-throughput sequencing, created a culture of large-scale biological and nutritional data. The expansion of big data and cloud computing in the 2000s further empowered the analysis of massive ingredient, supply chain, and consumer preference datasets. Since the beginning of the 21st century, advances in computational chemistry and cheminformatics in the 2010s have enabled the prediction of flavor molecules, aroma profiles, and chemical interactions. Breakthroughs in deep learning, exemplified by AlexNet in 2012, catalyzed applications in food recognition and image-based dietary tracking. Since 2014, generative models have expanded into food design, flavor and recipe generation, and ingredient discovery, with companies such as NotCo applying AI to plant-based food development. Since 2015, advanced fermentation and alternative protein models have started to include AI-driven strain engineering and process optimization to accelerate precision fermentation and industrial adoption. The current decade is witnessing the integration of wearable and microbiome data for AI-driven personalized nutrition since 2016, the emergence of digital twins for real-time monitoring and optimization of food manufacturing processes, and the rise of foundation models and large language models that leverage unstructured data for formulation and recipe innovation since the 2020s. Together, these advances are fueling the emergence of AI for Food as a distinct scientific discipline.
@@
Food is among the most powerful determinants of human health, planetary sustainability, and cultural identity [1]; yet, research and development for food remains slow, fragmented, and driven by trial-and-error rather than rational design. Radical, urgent, and creative food solutions are needed to combat the current climate crisis [2]. Notably, across farmland, livestock, and production, the food sector generates 35% of all global greenhouse gas emissions [3], and contributes heavily to biodiversity loss and public health challenges [4]. To meet climate goals and feed a growing population, we need alternative sources of protein that are delicious, affordable, scalable, and nutritionally compelling [5]. Today, however, the scientific bottlenecks in texture, flavor, scalability, and cost pose barriers to adoption [6]. Artificial intelligence (AI) has reshaped domains, from protein design and drug discovery to materials engineering [7]. AI is now poised to revitalize how we grow and make food–from optimizing plant-based meats to generating climate-conscious recipes [8]. Enabling protein diversification will have profound implications, and AI can substantially accelerate scientific progress if applied responsibly and towards the correct challenges [9]. But the success of AI is by no means guaranteed. When deployed in a scattershot way, AI can generate low-quality outputs, increase environmental burdens, or undermine consumer trust, aspects that are especially sensitive in food [10]. The goal is thus not to deploy AI indiscriminately. Instead, we need to partner with AI where it measurably reduces time, cost, and uncertainty for problems that matter [11]. Responsible application–focused on the right challenges–is essential. In this review, we identify areas across the food system where we can constructively engage with AI: we provide background on the current state of the art, outline new opportunities, and discuss the challenges that AI is facing in the quest to accelerate the future of food.
@@
The intellectual and technological trajectory that has shaped to the emergence of AI for Food as a discipline spans more than seven decades, weaving together advances in AI, food science, and biotechnology (Fig. 1). From the birth of symbolic AI in the 1950s [12], and the rise of machine learning and neural networks in the 1980s–90s [13], to the deep learning breakthroughs in the 2010s [14], and the advent of generative models [15] and foundation models [16] in the 2020s, AI has progressively expanded its ability to model, design, and predict complex systems. In parallel, developments in genomics [17], computational chemistry [18], fermentation, and personalized nutrition [19] have created unprecedented data resources and biological insights directly relevant to food design. Their convergence has given rise to a new field in which AI is now central to ingredient discovery, recipe formulation, manufacturing optimization, and consumer personalization [11].
@@
Designing food materials poses unique challenges, as their properties emerge from multiscale interactions among proteins, carbohydrates, lipids, and processing conditions [20]. Yet, AI is shifting food innovation from trial-and-error to data-driven discovery. Modern models link chemical composition to sensory outcomes and enable targeted formulation and rapid iteration. Machine-learning frameworks predict consumer preferences from chemical–sensory panels [21], while graph neural networks map molecular structure to odor quality [22]. Advances in protein modeling and molecular learning accelerate functional ingredient discovery [23, 24], and precision fermentation provides scalable routes to novel proteins and textures [25]. Together, these advances address long-standing challenges—engineering egg-free emulsions, meltable dairy alternatives, or fibrous plant-based meats—and have translated into market-ready products from companies such as NotCo, Perfect Day, and Climax Foods [26]. This transition from conceptual promise to industrial deployment underscores AI’s growing role as a driver of culinary innovation and sustainability.
We focus on the post-harvest stages and follow an AI-powered cycle through ingredient design, formulation development, fermentation and production, texture analysis, sensory properties, manufacturing, and recipe generation (Fig. 2). In the following sections, we highlight how AI-powered approaches, from generative models to digital twins, drive innovation across these steps.
@@
Results
@@
Ingredients: Driving functionality-first design. Novel food ingredients must balance sustainability, nutrition, sensory appeal, and functionality. Functionality includes water-holding, emulsification, gelation, viscosity, and stability that shape processing and structure. Many ingredients aim to replicate animal-derived functions, such as foaming in egg replacers; yet, matching performance remains challenging. Structure–function relationships link molecular architecture to activity. In food, macromolecular features of proteins, lipids, and fibers encode functionality. This raises the question to which extent AI can bridge this scale gap and translate molecular features into techno-functional predictions for a functionality-first design–build–test–learn cycle.
The success of animal-free ingredients depends on functionality and the ability to meet formulation and consumer requirements [28]. These traits arise from molecular features, environmental conditions such as pH and temperature, and interactions with fats, fibers, and solutes. Proteins, lipids, and fibers provide core functionality, with proteins most versatile. The most promising alternative proteins today are wheat gluten and soy; they can form fibrous textures under extrusion and shear [28]. New sources include legumes, oilseeds, grains, algae, and leaves [29]. Processing strongly shapes functionality: dry fractionation preserves native structures, while wet fractionation produces isolates that control solubility and gelation [30]. Lipids and fibers add complementary features, from stabilizing interfaces to modulating viscosity and lubrication. Since plant heterogeneity limits predictability, developers modify ingredients chemically, physically, or enzymatically. The field is now shifting toward functionality-enriched fractions for more consistent, cost-effective performance [29].
@@
AI and informatics enable functionality-driven ingredient design by linking molecular and mesoscopic features to macroscopic outcomes under processing constraints. Classification models support discovery; for example, a QSAR model identified glycyrrhizin as a natural emulsifier [31]. Regression models predict properties such as pectin viscosity to reduce cost and accelerate formulation [32]. Machine learning can predict techno-functional traits across crops with ${\textsf{R}}^{2}>0.80$ [33], while AI models can capture plant protein rheology with ${\textsf{R}}^{2}>0.99$ [34]. Deep learning increasingly links functionality to sensory design [35]. A recent study optimized starch-based gels for saltiness and strength using e-tongue and oral-processing data [36]. Multi-objective optimization can predict solubility, gel strength, and emulsification towards scalable formulation strategies [33]. Together, these advances support closed-loop design–build–test–learn cycles. High-throughput screening, active learning, and digital twins operationalize models, while generative approaches target multiple criteria and bridge lab-to-plate gaps.
@@
Ingredient design still struggles with consistent performance across food matrices and key challenges remain. First, models often fail to generalize across ingredients and conditions. Second, nonlinear interactions in multicomponent systems complicate prediction and require hybrid physics-informed approaches. Third, sensory preferences vary across cultures, which demands diverse datasets. Fourth, models must link molecular features such as sequence, charge, and hydrophobicity to functionality in a transparent and scalable way. Addressing these challenges requires open, collaborative tools and an AI-driven ecosystem that predicts techno-functional performance from molecular features.
@@
Formulations: Mapping flavor and recipe spaces. Ingredient functionality sets the foundation, but effective food design depends on how those ingredients interact in complex systems; formulation brings these components together. Meeting the dual mandate of sustainability and sensory fidelity requires moving beyond slow trial-and-error formulation development. AI offers a compelling approach to uncover hidden patterns that connect ingredients, formulations, production, texture, and sensory properties, and propose candidate formulations that are predicted to meet nutritional, functional, or sensorial targets [37, 38, 39]. Today, AI practitioners can draw on a broad collection of public resources including Recipe1M+ [40], FlavorGraph [41], FlavorDB [42], USDA FoodData Central, and FAO/INFOODS, as well as domain-specific databases such as Phenol-Explorer, LipidBank, and the Volatile Compounds in Food database. Inputs include molecular descriptors, nutritional and functional properties, recipe structures, and process sequences; outputs are complete candidate formulations aligned with target criteria. AI-powered approaches combine latent-space learning through autoencoders with conditional generation, constraint-aware optimization, and expert-in-the-loop refinement (Fig. 3A). But can they truly capture the complex, non-linear interactions between ingredients, processes, and perception–even under sparse and noisy data–and unlock a new generation of food formulations?
Figure 3: AI for food. (A) AI for Formulations. An iterative workflow to discover new formulations encodes target profiles such as sensory properties, cost, and nutritional profiles; generates formulation sets with deep learning; predicts properties and scores; and integrates laboratory prototyping and sensory panels in continuous feedback loops. (B) AI for Texture. Machine learning integrates multimodal data, from ingredients to process conditions, to optimize plant-based meat analogues, from microstructural appearance to macrostructural texture. Future machine learning tools could help reverse engineer ingredients and process conditions to achieve a desired appearance and texture. (C) AI for Fermentation and Production. Artificial intelligence and large language models can accelerate the fermentation and production cycle during microbial strain selection, data collection, process optimization, real time monitoring and adaptive control, upscaling of prediction and simulation, and product profiling and customization. (D) AI for Sensory Properties. Optimizing the sensory perception of food is traditionally a costly, time-consuming process involving expert tasting panels and instrumental analysis, including mass spectrometry and gas chromatography. In contrast, an AI-powered pipeline has the potential to predict sensory perceptions from only the chemical or molecular make-up of food, eliminating the need for costly human panels and reducing the time to iterate product formulations.
@@
Computational gastronomy has long shown that recipes and flavor chemistry form structured networks that we can mine for predictive patterns. The Flavor Network formalized ingredient compatibility via shared volatiles [43], while FlavorGraph integrated one million recipes with $\sim$1,500 flavor molecules to generate graph embeddings for pairing prediction and ingredient clustering [41]. These studies have successfully shown that high-dimensional, multimodal food representations can capture the complexity of culinary design. In alternative proteins, industrial disclosures from NotCo patents [37, 38, 39] illustrate hybrid generative optimization pipelines. These systems embed ingredients and recipes in latent spaces using autoencoders and recurrent networks, condition sampling on feature vectors of target product profiles, refine candidates with constraint solvers such as sparse regression, and iteratively incorporate expert sensory feedback. This architecture mirrors multimodal generative trends in other fields and enables analogue discovery. It can identify non-obvious ingredient substitutions, while maintaining process feasibility. Other companies, including Climax Foods and Eat Just, deploy similar proprietary approaches, while academic research continues to build on Recipe1M+ [40], FlavorGraph [41], and FlavorDB [42]. Together, these efforts highlight a state of the art that blends representation learning with constraint-aware generation to accelerate formulation discovery.
@@
Generative AI expands opportunities across protein diversification. For example, it can design plant-based cheeses that melt and stretch by identifying proteins that act like casein, the key driver of structure and meltability in animal cheese. It can discover ingredient blends that replicate meat textures and pinpoint volatile combinations that evoke familiar flavors [44]. And it can improve for plant-based meats from sensory panel feedback and match expert predictions of panel rankings [45]. Large-scale screening lets scientists focus on a few promising prototypes instead of hundreds of trials, which shortens the development cycle. In the near future, these tools will be able to tailor formulations to individual nutritional needs and taste preferences and enable personalized food design.
@@
But significant challenges remain before AI can transform formulation discovery at scale. High-quality sensory datasets are limited and often fail to generalize across different food types. Even with detailed chemical and nutritional data, predicting human perception remains an open problem. Modeling ingredient performance under varying heat, shear, or storage conditions is incomplete. Cost, supply chain, allergen control, and clean-label requirements must be built into models as hard constraints. Trust among regulators, industry, and consumers will depend on transparent, auditable design processes. The near-term path will be hybrid, where AI proposes and humans refine, toward transparent formulation platforms that integrate constraints and predict performance at scale.
Fermentation: Enabling smart bioreactors. Molecular and formulation design define the building blocks, but realizing performance requires precise control at the cellular and microbial scale. Precision fermentation—where engineered microorganisms are programmed to produce functional proteins such as casein or heme—has become central to protein diversification and sustainable food production. Its success depends on tightly coupled variables, strain stability, metabolic fluxes, nutrient composition, and bioreactor dynamics. AI has the potential to fuse these heterogeneous data with mechanistic insight to enable predictive, scalable, and cost-effective fermentation and production to address critical bottlenecks such as yield optimization and resource efficiency (Fig. 3C).
@@
Optimizing fermentation requires a deep understanding of bioprocesses to maximize yield and efficiency while maintaining homogeneity, sterility, and scalability [46]. Kinetic bioprocess models remain widely used but are constrained by steady-state assumptions, and metabolic flux models for engineered strains often face computational bottlenecks. Experimental optimization through orthogonal design [47] and response surface methodology [48] remains the industry standard, while multiscale feature engineering is still limited [49]. Machine learning has proven valuable for modeling nonlinear fermentation dynamics [50] and for real-time monitoring of anaerobic digestion [51]. For example, deep learning soft sensors that combine convolutional neural networks with bidirectional long short-term memory architectures have achieved ${\textsf{R}}^{2}$ values up to 0.98 in predicting key anaerobic digestion state variables from real-time pH, electrical conductivity, and oxidation–reduction potential measurements [51]. While AI and reinforcement learning have been successfully applied to sustainable protein fermentation [52] and co-culture optimization [53], large-scale adoption of AI, and especially generative AI, remains limited.
@@
AI can shift fermentation and production from reactive optimization to predictive management, with customized strains and self-optimizing processes. Machine learning models can optimize strain design, nutrient feeds, and bioreactor control, and deep learning approaches can capture non-linear interactions better than traditional regression [54]. Automation and AI-enhanced imaging [55], soft sensors [56], and spectroscopy [52] can improve monitoring and adaptive control. Algorithms such as support vector machines [57] and fuzzy inference systems [58] support robust predictions, while evolutionary optimization [59] aids process tuning. Hybrid mechanistic-AI models [60] and explainable AI frameworks [61] can improve transparency and decision-making. Training LLMs on unstructured bioprocess data–including text, images, and sensor logs–could support knowledge retrieval, troubleshooting, and experimental design and bridge biology, chemistry, engineering, and design through actionable insights, like predicting protein structures or cell morphology [62].
@@
To realize this potential, the field must overcome the key barriers of data scarcity, data integration, and model transferability. Generating large experimental datasets remains labor intensive, but scaled-down bioreactors and small-data algorithms now offer new possibilities. At the same time, robust biosecurity and ethical frameworks will be critical to safely employ generative AI. Integration with genome-scale metabolic models and digital twins of bioreactor performance could create predictive platforms for design and scale-up. The collective goal is an integrated fermentation platform that links strain design, bioreactor dynamics, and real-time sensing to support predictive scale-up and reliable process optimization.